The sum of the first eight terms of an A.P. is 4 and the sum of the first twelve terms is 78. Find the sum of the first 30 terms.
Answer by stanbon(75887)

The sum of the first eight terms of an A.P. is 4 and the sum of the first twelve terms is 78. Find the sum of the first 30 terms.
--------
1st 8 terms:: 4[a(1)+a(1)+3d] = 8a(1) + 12d = 4
--------
1st 12 terms: 6[a(1)+a(1)+11d] = 12a(1)+ 66d = 78
--------
Multiply thru the 1st equation by (3/2) to get::
12a(1) + 18d = 6
12a(1) + 66d = 78
-------------
48d = 72
d = 3/2
----
Solve for "a(1)"::
8a(1) + 18d = 4
8a(1) + 27 = 4
-----
8a(1) = -23
a(1) = -23/8
-------
Ans: S(30) = 15[(-23/8) + [-23/8+19(3/2)]] =
Do the arithmetic.
Cheers,
Stan H.
-------------